The Franklin County health department is investigating vaccination rates and effectiveness for the HPV vaccine among college students. 424 college students from area colleges were studies. Vaccination status and infection status was obtained by medical record. Results are shown in the table below:

Active Infection Recovered Infection Not Infected
Vaccinated 22 30 176
Not Vaccinated 46 48 102

Suppose a randomly selected student is not infected. What is the probability that this student has been vaccinated?Calculate the probability that a randomly selected college student from Franklin County has been vaccinated.

Calculate the probability that a randomly selected college student from Franklin County has an active infection.

The health commissioner is interested in reducing the rates of active infection. Describe what would happen to the probability of active infection if vaccination rates increased in Franklin County. Provide specific numbers or calculations from the question to support your answer.
Answer #1
Active infection recovered infection not infected total
vaccinated 22 30 176 228
not vaccinated 46 48 102 196
total 68 78 278 424

A)
P(V | NI) = P(V and NI)/P(NI) = 176/278 = 0.6331

B)
P(vaccinated) = 228/278 = 0.8201

C)
P(AI or RI | NV) = (48 + 102)/196 = 0.7653

D)
Active infection in case of vaccination is very low.
hence increasing the vaccination reduces the active infection